OpenStudy (anonymous):

WORD PROBLEM: 2 variable & 2 equations. Let X = miles of 1st car traveled. Let Y = miles of 2nd car traveled. "A family has two cars. The first car has a fuel efficiency of 20 miles per gallon of gas and the second has a fuel efficiency of 40 miles per gallon of gas. During one particular week, the two cars went a combined total of 1400 miles, for a total gas consumption of 55 gallons. How many gallons were consumed by each of the two cars that week?"

OpenStudy (anonymous):

Ok, so there is a statement about total number of miles and a statement about total number of gallons. Those will make your two equations.

OpenStudy (anonymous):

First equation I got is: x + y = 1400

OpenStudy (anonymous):

I'm not sure what eq2 is

OpenStudy (anonymous):

Good. Note that miles ÷ miles/gallon = number of gallons.

OpenStudy (anonymous):

OK. I'm trying to figure out Eq2 but having a hard time. Give me another moment

OpenStudy (anonymous):

is it 20x + 40y =55 ?

OpenStudy (anonymous):

or is it x/20 + y/40 = 55

OpenStudy (anonymous):

The second one.

OpenStudy (anonymous):

Ok let me try to figure this out then

OpenStudy (anonymous):

It's a pretty easy one once you clear the fractions. You can probably do it mentally.

OpenStudy (anonymous):

x = 800 and y = 600. I checked it in 2nd equation and got 55 gal.